adovbush:

The volume of a cube can be found using the equation V = s³, where V is the volume and s is the measure of one side of the cube. Which equation can be used to solved to find the side length of a cube with a volume of 999 in³?

adovbush:

s=999−−−√3 in. s=9993 in. s=999 in. s=999−−−√2 in.

adovbush:

The 3 lines are on top of the 999's

adovbush:

@Shadow

Shadow:

Draw it, not sure what you're referring to

adovbush:

1 attachment
Shadow:

What do you think the answer is?

adovbush:

A or D

Shadow:

Hmm, why those?

adovbush:

I'm just guessing cause I have no idea how to solve it

Shadow:

Lets say we have a cube. Remember that all sides are the same length. One side is 5 ft long. If I want to solve for the volume, I need to solve for the area (s times s) multiplied by the height, (also 5). So we have: \[5 \times 5 \times 5 = 5^3\] \[5^3 = 125ft\]

Shadow:

125ft is the volume. If I were to want to get the sides again, I would take the 3rd root of 125. Also written as: \[\sqrt[3]{125} = 5\]

Shadow:

S = 5

Shadow:

\[S = \sqrt[3]{125}\]

Shadow:

Does that example make sense to you?

adovbush:

Kind of but does that mean i have to do 125 x 125 x 125?

adovbush:

Since S = \[3\sqrt{}\]125

adovbush:

Woops well u know what i mean

Shadow:



Shadow:

Taking the 3rd root of 125 is not 125 times 125 times 125. It is you finding breaking 125 down to find three common factors.

Shadow:

Example:\[\sqrt[3]{81} = \sqrt[3]{9 \times 9} = \sqrt[3]{3 \times 3 \times 3 \times 3} = 3\sqrt[3]{3}\]

Shadow:

\[\sqrt[3]{8} = \sqrt[3]{2 \times 2 \times 2} = 2\]

Shadow:

Is this making sense?

adovbush:

Nope not even a bit.

Shadow:

What part are you not getting?

adovbush:

I have no idea what this thing is : \[\sqrt{}\]

adovbush:

Shadow, thank you so much for your help, even though I don't get it. I will ask my mom right now, and she will help me. I'm so sorry for taking away so much time from you :(

Shadow:

All good man. Good luck

adovbush:

Thanks))